# Repurposing Synthetic Data for Fine-grained Search Agent Supervision

### v1

@@ Abstract
LLM-based search agents are increasingly trained on entity-centric synthetic data to solve complex, knowledge-intensive tasks. However, prevailing training methods like Group Relative Policy Optimization (GRPO) discard this rich entity information, relying instead on sparse, outcome-based rewards. This critical limitation renders them unable to distinguish informative "near-miss" samples—those with substantially correct reasoning but a flawed final answer—from complete failures, thus discarding valuable learning signals. We address this by leveraging the very entities discarded during training. Our empirical analysis reveals a strong positive correlation between the number of ground-truth entities identified during an agent’s reasoning process and final answer accuracy. Building on this insight, we introduce Entity-aware Group Relative Policy Optimization (E-GRPO), a novel framework that formulates a dense entity-aware reward function. E-GRPO assigns partial rewards to incorrect samples proportional to their entity match rate, enabling the model to effectively learn from these "near-misses". Experiments on diverse question-answering (QA) and deep research benchmarks show that E-GRPO consistently and significantly outperforms the GRPO baseline. Furthermore, our analysis reveals that E-GRPO not only achieves superior accuracy but also induces more efficient reasoning policies that require fewer tool calls, demonstrating a more effective and sample-efficient approach to aligning search agents.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has catalyzed the development of sophisticated autonomous agents, with search agents emerging as a prominent class for solving complex, knowledge-intensive tasks (Yao et al., 2023; Wang et al., 2024; Xi et al., 2025). Training these agents to navigate the vast, noisy web effectively requires abundant and challenging data (Google Team, 2025; OpenAI, 2025; xAI Team, 2025; Moonshot AI, 2025). To meet this demand, a dominant paradigm of synthetic data generation has emerged (Wu et al., 2025b; Li et al., 2025b; Gao et al., 2025). In this paradigm, as shown in Figure 1 (left), complex questions are often created by systematically transforming simple "seed" questions through operations like fact injection or deliberate obfuscation. This process creates an intricate problem structure, paved with key entities that form the factual backbone of the correct answer.
This synthetic data is then used to train agents within the now-dominant reinforcement learning (Wen et al., 2024; Singh et al., 2025), especially with Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and its numerous variants (Yu et al., 2025; Dong et al., 2025; Hu, 2025; Xu et al., 2024; Zhao et al., 2025). These methods typically rely on outcome-based rewards, utilizing only the final answer while discarding the intermediate entity information meticulously embedded during data synthesis. This mechanism leads to the reward sparsity problem (Qian et al., 2025; Deng et al., 2025), which manifests critically for search agents (Song et al., 2025; Wu et al., 2025a; Jin et al., 2025; Li et al., 2025c; Zheng et al., 2025; Zhang et al., 2025a; Li et al., 2025b; Gao et al., 2025): by treating all negative samples uniformly, GRPO fails to distinguish a "near-miss"—a response with correct intermediate reasoning steps but a flawed answer—from a complete failure. For instance, in answering Who was the director of the 1997 film starring the actor who won the Academy Award for Best Actor for the film ’The Revenant’?, a "near-miss" that correctly identifies the actor (Leonardo) and the film (Titanic) but fails on the final answer is far more informative than one that misunderstands the query entirely. By penalizing both equally, standard GRPO discards crucial learning signals embedded in partially correct reasoning, forcing the model to re-learn steps it had already mastered.
@@
Based on this core insight, we propose Entity-aware Group Relative Policy Optimization (E-GRPO), a novel RL framework that enhances policy optimization by formulating a dense, entity-aware reward function from the entities within the synthetic training data. Specifically, instead of applying a uniform penalty, our method assigns a bonus to negative samples proportional to their entity match rate. By doing so, a "near-miss" sample, which contains many correct entities and is highly informative for learning, receives a better reward than a complete failure. This fine-grained reward, obtained with negligible computational cost, compels the model to move beyond simply avoiding errors and towards mastering the process of identifying and synthesizing key information, thereby addressing the limitation of standard GRPO in complex search tasks.
@@
We identify the "near-miss" problem in GRPO-based training and propose the core insight that entities from synthetic data can be repurposed as a fine-grained reward signal, supported by empirical analysis revealing a strong correlation between entity match rate and task accuracy.
@@ 3.1 Analyzing Entity Matching in Agentic Reasoning
An entity is considered "mentioned" if its full phrase appears as an exact string match in the thought’s text (more discussion available in Appendix B). The entity match rate for rollout , denoted as , is then calculated as the ratio of matched entities to the total:
@@
To investigate the correlation between entity match rate and accuracy, we first conduct a per-question analysis on a sampled subset of SailorFog-QA (Li et al., 2025b) using the WebSailor-7B agent (Li et al., 2025b). For each question, we perform 8 rollouts and calculate the average entity match rates of correctly solved and failed rollouts, respectively. As shown in Figure 1 (upper right), the average entity match rate of correct rollouts was higher than that of failed ones in the vast majority of the questions, outnumbering the reverse scenario by a clear 4-to-1 margin (1939 vs. 487 questions). This establishes a strong correlation between the entity match rate and the correctness of the final answer.
Moving beyond this aggregate, per-question view, we analyze the distribution of the normalized entity match rate across all individual rollouts. As shown in Figure 1 (bottom right), the distributions for correct and incorrect rollouts diverge significantly. The distribution of correct samples (green) peaks sharply at a normalized rate of 1.0. In contrast, incorrect samples (red) show a bimodal distribution: a large peak at 0.0, and a notable spread across the mid-to-high range. This latter group represents the informative "near-misses", where most entities were found but the final reasoning failed.
This analysis shows that the entity match rate is more than just a pass/fail indicator. Instead, it provides a granular scale to distinguish valuable "near-misses" from complete failures, offering a richer signal of an agent’s reasoning quality.
@@ 3.2 Entity-aware Group Relative Policy Optimization
The limitation of this formulation is evident: as shown in Figure 2, standard GRPO assigns an identical reward of 0 to both a complete failure (middle rollout, 0 entities matched) and an informative "near-miss" (bottom rollout, 1 entity matched), thus rendering their different reasoning qualities indistinguishable.
@@
where is a hyperparameter balancing the value of accuracy and entity matching. This formulation yields two significant advantages. (1) It creates a dense reward spectrum to distinguish the quality of incorrect rollouts. As shown in Figure 2, a "near-miss" that identifies a correct entity (Leonardo) is rewarded (), unlike a complete failure which receives zero. (2) It provides a meaningful training signal even in all-wrong groups where standard GRPO offers no gradient.
@@
KL-Free Objective and Policy Exploration. Following DAPO (Yu et al., 2025), we remove the KL-divergence regularization term in GRPO and apply the "clip-higher" method, which increases the upper clipping bound , to better encourage policy exploration.
@@ 4.1 Experiment Setup
RL: We generate two distinct 1k-sample datasets for RL training. For the Local environment, we synthesize data using the Asearcher (Gao et al., 2025) method over the 2024 Wikipedia corpus. For the Web environment, we use the SailorFog-QA data generation pipeline. Crucially, for both datasets, we retain all ground-truth entities generated during the synthesis process to enable E-GRPO. We train the 7B model in both environments, while the 30B model is trained only in the more complex Web environment. For each setup, we apply both GRPO and E-GRPO.
@@
| Comparison with Other Baselines | |||||||||
@@ 4.3 Analysis
We begin by analyzing the training dynamics of E-GRPO against the GRPO baseline. As shown in Figure 3, E-GRPO demonstrates superior learning efficiency and effectiveness. The left panel shows that E-GRPO (purple) consistently achieves higher training accuracy, showing a steadier and more pronounced upward trend than the GRPO baseline (blue). This suggests that the dense, entity-aware reward provides a more effective and stable learning signal. Simultaneously, the middle panel reveals that E-GRPO learns a more efficient reasoning policy, consistently using fewer tool calls per rollout. This efficiency can be attributed to rewarding the discovery of key entities, which guides the agent towards more direct and informative solution steps.
@@
A detailed case study in Appendix D provides a qualitative illustration of these dynamics, concretely demonstrating how E-GRPO’s focus on entity matching leads to a more efficient and accurate reasoning path.
@@
We conduct an ablation study on the hyperparameter , which balances the outcome-based reward and the entity-matching bonus. As shown in Figure 4, setting reduces our method to the GRPO baseline. For all four benchmarks, performance consistently improves as increases from 0.0, peaking at 0.3. This clearly demonstrates the benefit of incorporating the entity-aware reward. However, a further increase to leads to a performance drop on most benchmarks, suggesting that an excessively strong entity-matching bonus can distract the model from the primary goal of generating a correct final answer. This highlights the importance of balancing the two reward components, with a moderate value yielding the optimal policy.
@@ 5 Related Work
The capabilities of Large Language Models (LLMs) have fueled a surge in research on autonomous agents that can interact with external environments to solve complex tasks. A foundational paradigm in this area is the ReAct framework (Yao et al., 2023), which interleaves reasoning (thought) and action steps. Building on this, a prominent line of research has focused on search agents designed to navigate the web (Song et al., 2025; Zheng et al., 2025; Li et al., 2025c; Zhang et al., 2025a; Sun et al., 2025). Advanced models like Gemini Deep Research (Google Team, 2025), OpenAI Deep Research (OpenAI, 2025), Grok DeepSearch (xAI Team, 2025), along with smaller open-source models such as Asearcher (Gao et al., 2025), WebThinker (Li et al., 2025d), WebWalker (Wu et al., 2025b), WebDancer (Wu et al., 2025a), and WebSailor (Li et al., 2025b; a) have demonstrated increasing proficiency in retrieving and synthesizing information from noisy, real-world web sources. Our work directly contributes to improving the training methodology for this class of agents, addressing the challenge of learning robust policies in complex web environments.

### v2

@@ Abstract
LLM-based search agents are increasingly trained on entity-centric synthetic data to solve complex, knowledge-intensive tasks. However, prevailing training methods like Group Relative Policy Optimization (GRPO) discard this rich entity information, relying instead on sparse, outcome-based rewards. This critical limitation renders them unable to distinguish informative “near-miss” samples—those with substantially correct reasoning but a flawed final answer—from complete failures, thus discarding valuable learning signals. We address this by leveraging the very entities discarded during training. Our empirical analysis reveals a strong positive correlation between the number of ground-truth entities identified during an agent’s reasoning process and final answer accuracy. Building on this insight, we introduce Entity-aware Group Relative Policy Optimization (E-GRPO), a novel framework that formulates a dense entity-aware reward function. E-GRPO assigns partial rewards to incorrect samples proportional to their entity match rate, enabling the model to effectively learn from these “near-misses”. Experiments on diverse question-answering (QA) and deep research benchmarks show that E-GRPO consistently and significantly outperforms the GRPO baseline. Furthermore, our analysis reveals that E-GRPO not only achieves superior accuracy but also induces more efficient reasoning policies that require fewer tool calls, demonstrating a more effective and sample-efficient approach to aligning search agents.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has catalyzed the development of sophisticated autonomous agents, with search agents emerging as a prominent class for solving complex, knowledge-intensive tasks (Yao et al., 2023; Wang et al., 2024; Xi et al., 2025). Training these agents to navigate the vast, noisy web effectively requires abundant and challenging data (Google Team, 2025b; OpenAI, 2025; xAI Team, 2025; Moonshot AI, 2025). To meet this demand, a dominant paradigm of synthetic data generation has emerged (Wu et al., 2025b; Li et al., 2025b; Gao et al., 2025). In this paradigm, as shown in Figure 1 (left), complex questions are often created by systematically transforming simple “seed” questions through operations like fact injection or deliberate obfuscation. This process creates an intricate problem structure, paved with key entities that form the factual backbone of the correct answer.
This synthetic data is then used to train agents within the now-dominant reinforcement learning (Wen et al., 2024; Singh et al., 2025), especially with Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and its numerous variants (Yu et al., 2025; Dong et al., 2025; Hu, 2025; Xu et al., 2024; Zhao et al., 2025). These methods typically rely on outcome-based rewards, utilizing only the final answer while discarding the intermediate entity information meticulously embedded during data synthesis. This mechanism leads to the reward sparsity problem (Qian et al., 2025; Deng et al., 2025), which manifests critically for search agents (Song et al., 2025; Wu et al., 2025a; Jin et al., 2025; Li et al., 2025c; Zheng et al., 2025; Zhang et al., 2025a; Li et al., 2025b; Gao et al., 2025): by treating all negative samples uniformly, GRPO fails to distinguish a “near-miss”—a response with correct intermediate reasoning steps but a flawed answer—from a complete failure. For instance, in answering Who was the director of the 1997 film starring the actor who won the Academy Award for Best Actor for the film ’The Revenant’?, a “near-miss” that correctly identifies the actor (Leonardo) and the film (Titanic) but fails on the final answer is far more informative than one that misunderstands the query entirely. By penalizing both equally, standard GRPO discards crucial learning signals embedded in partially correct reasoning, forcing the model to re-learn steps it had already mastered.
@@
Based on this core insight, we propose Entity-aware Group Relative Policy Optimization (E-GRPO), a novel RL framework that enhances policy optimization by formulating a dense, entity-aware reward function from the entities within the synthetic training data. Specifically, instead of applying a uniform penalty, our method assigns a bonus to negative samples proportional to their entity match rate. By doing so, a “near-miss” sample, which contains many correct entities and is highly informative for learning, receives a better reward than a complete failure. This fine-grained reward, obtained with negligible computational cost, compels the model to move beyond simply avoiding errors and towards mastering the process of identifying and synthesizing key information, thereby addressing the limitation of standard GRPO in complex search tasks.
@@
We identify the “near-miss” problem in GRPO-based training and propose the core insight that entities from synthetic data can be repurposed as a fine-grained reward signal, supported by empirical analysis revealing a strong correlation between entity match rate and task accuracy.
@@ 3.1 Analyzing Entity Matching in Agentic Reasoning
An entity is considered “mentioned” if its full phrase appears as an exact string match in the thought’s text (more discussion available in Appendix B). The entity match rate for rollout , denoted as , is then calculated as the ratio of matched entities to the total:
@@
To investigate the correlation between entity match rate and accuracy, we first conduct a per-question analysis on a sampled subset of SailorFog-QA (Li et al., 2025b) using the WebSailor-7B agent (Li et al., 2025b). For each question, we perform 8 rollouts and calculate the average entity match rates of correctly solved and failed rollouts, respectively (further explanation in Appendix I). As shown in Figure 1 (upper right), the average entity match rate of correct rollouts was higher than that of failed ones in the vast majority of the questions, outnumbering the reverse scenario by a clear 4-to-1 margin (1939 vs. 487 questions). This establishes a strong correlation between the entity match rate and the correctness of the final answer.
Moving beyond this aggregate, per-question view, we analyze the distribution of the normalized entity match rate across all individual rollouts. As shown in Figure 1 (bottom right), the distributions for correct and incorrect rollouts diverge significantly. The distribution of correct samples (green) peaks sharply at a normalized rate of 1.0. In contrast, incorrect samples (red) show a bimodal distribution: a large peak at 0.0, and a notable spread across the mid-to-high range. This latter group represents the informative “near-misses”, where most entities were found but the final reasoning failed.
This analysis shows that the entity match rate is more than just a pass/fail indicator. Instead, it provides a granular scale to distinguish valuable “near-misses” from complete failures, offering a richer signal of an agent’s reasoning quality.
@@ 3.2 Entity-aware Group Relative Policy Optimization
The limitation of this formulation is evident: as shown in Figure 2, standard GRPO assigns an identical reward of 0 to both a complete failure (middle rollout, 0 entities matched) and an informative “near-miss” (bottom rollout, 1 entity matched), thus rendering their different reasoning qualities indistinguishable.
@@
where is a hyperparameter balancing the value of accuracy and entity matching. This formulation yields two significant advantages. (1) It creates a dense reward spectrum to distinguish the quality of incorrect rollouts. As shown in Figure 2, a “near-miss” that identifies a correct entity (Leonardo) is rewarded (), unlike a complete failure which receives zero. (2) It provides a meaningful training signal even in all-wrong groups where standard GRPO offers no gradient.
@@
KL-Free Objective and Policy Exploration. Following DAPO (Yu et al., 2025), we remove the KL-divergence regularization term in GRPO and apply the “clip-higher” method, which increases the upper clipping bound , to better encourage policy exploration.
@@ 4.1 Experiment Setup
RL: We generate two distinct 1k-sample datasets for RL training. For the Local environment, we synthesize data using the Asearcher (Gao et al., 2025) method over the 2024 Wikipedia corpus. For the Web environment, we use the SailorFog-QA data generation pipeline. Note that both methods are anchored in entities from Wikipedia despite the distinct environments they use. Crucially, for both datasets, we retain all ground-truth entities generated during the synthesis process to enable E-GRPO. We train the 7B model in both environments, while the 30B model is trained only in the more complex Web environment. For each setup, we apply both GRPO and E-GRPO.
@@
| Comparison with Other Reference Agents | |||||||||
@@ 4.3 Analysis
We begin by analyzing the training dynamics of E-GRPO against the GRPO baseline. As shown in Figure 3, E-GRPO demonstrates superior learning efficiency and effectiveness. The left panel shows that E-GRPO (purple) consistently achieves higher training accuracy, showing a steadier and more pronounced upward trend than the GRPO baseline (blue). This suggests that the dense, entity-aware reward provides a more effective and stable learning signal. Simultaneously, the middle panel reveals that E-GRPO learns a more efficient reasoning policy, consistently using fewer tool calls per rollout. This efficiency can be attributed to rewarding the discovery of key entities, which guides the agent towards more direct and informative solution steps. Extended training dynamics are provided in Appendix G for reference.
@@
A detailed case study in Appendix D provides a qualitative illustration of these dynamics, concretely demonstrating how E-GRPO’s focus on entity matching leads to a more efficient and accurate reasoning path. Through the case study, we also analyze several failure cases of E-GRPO in Appendix E.
@@
We conduct an ablation study on the hyperparameter , which balances the outcome-based reward and the entity-matching bonus. As shown in Figure 4, setting reduces our method to the GRPO baseline. For all four benchmarks, performance consistently improves as increases from 0.0, peaking at 0.3. This clearly demonstrates the benefit of incorporating the entity-aware reward. However, a further increase to leads to a performance drop on most benchmarks, suggesting that an excessively strong entity-matching bonus can distract the model from the primary goal of generating a correct final answer. This highlights the importance of balancing the two reward components, with a moderate value yielding the optimal policy. More analysis of a decaying value during training is available in Appendix F.
@@ 5 Related Work
The capabilities of Large Language Models (LLMs) have fueled a surge in research on autonomous agents that can interact with external environments to solve complex tasks. A foundational paradigm in this area is the ReAct framework (Yao et al., 2023), which interleaves reasoning (thought) and action steps. Building on this, a prominent line of research has focused on search agents designed to navigate the web (Song et al., 2025; Zheng et al., 2025; Li et al., 2025c; Zhang et al., 2025a; Sun et al., 2025). Advanced models like Gemini Deep Research (Google Team, 2025b), OpenAI Deep Research (OpenAI, 2025), Grok DeepSearch (xAI Team, 2025), along with smaller open-source models such as Asearcher (Gao et al., 2025), WebThinker (Li et al., 2025d), WebWalker (Wu et al., 2025b), WebDancer (Wu et al., 2025a), and WebSailor (Li et al., 2025b; a) have demonstrated increasing proficiency in retrieving and synthesizing information from noisy, real-world web sources. Our work directly contributes to improving the training methodology for this class of agents, addressing the challenge of learning robust policies in complex web environments.
@@ 6 Conclusion
7 Acknowledgment
This work was supported by Alibaba Group through Alibaba Innovative Research Program, the Core Facility Platform of Computer Science and Communication, SIST, and the HPC platform of ShanghaiTech University.
@@ Appendix D Case Study
Appendix E Failure Analysis
Our extensive case study reveals that, in addition to the factual errors our paper targets at, agent failures frequently stem from several other distinct error modes, which we detail below.
Broken Reasoning Coherence: The agent loses reasoning coherence by failing to act on its previously stated goal, particularly after a Visit tool call.
Example:
In this example, the agent exhibits a loss of reasoning coherence. While it correctly visits the page and extracts a key fact (“a single frame”), it subsequently ignores its own goal for that visit (“notable architectural features”) and instead begins reasoning about an unrelated subject.
Distracted Querying: The agent loses focus on the core objective and instead pursues secondary information triggered by minor details, deviating from the optimal solution path, finally exceeding the context length.
Example:
In this example, the agent fails to maintain its core objective. It correctly formulates an early query that includes the temporal aspect (“early 1970s”), but then gets distracted by other entities (“MoMA”). The agent forgets that its primary task was to find the year, as shown by the absence of any time-related keywords in its later search query.
Information Overload from Concurrent Queries: By executing multiple search queries within a single step, the agent generates a large volume of unstructured text. This information overload overwhelms its processing capacity, causing it to fail to identify or retain the most critical facts necessary for the subsequent reasoning step.
Example:
In this example, the agent executes three highly similar queries simultaneously. While only the third query is well-targeted, containing the critical term “Chair”, the valuable information it returns is buried within the larger volume of text from all three searches. Overwhelmed by the dense context, the agent fails to isolate the key signal from the noise, which misleads it down a flawed reasoning path.
Our case study indicates E-GRPO mitigates broken reasoning coherence. By incentivizing the agent to carry forward critical entities into its thought process, the entity-aware reward implicitly enforces topical continuity and prevents the reasoning from drifting into irrelevance.
The other identified failure modes, however, suggest promising avenues for future work:
For distracted querying, a potential solution is to periodically re-inject the original question into the agent’s context, orienting the agent’s focus towards the primary objective.
We attribute information overload primarily to the inherent limitations of the base model in processing long, unstructured contexts. Addressing this could involve either leveraging more advanced models with superior long-context capabilities or designing a penalty mechanism for imprecise or redundant queries.
Appendix F Decaying Entity Matching Weights
| alpha | GAIA | BrowseComp | BrowseComp-ZH | xbench-DS |
| 0.3 | 48.5 | 12.9 | 26.4 | 46.7 |
| decay | 48.2 | 12.8 | 26.2 | 47.3 |
We also train the 30B model with an alpha that linearly decays from 0.3 to 0.0 over the first 60 steps (80 steps in total). The results are presented in Table 3.
Comparing with a fixed 0.3, the results show no clear or consistent advantage for the decaying alpha strategy. This suggests our paper’s choice of a simple, fixed alpha as a practical and effective setting. However, we believe a dynamic alpha holds potential. It’s possible that applying the decaying schedule over longer training horizons could be more impactful, which is a promising direction for future research.
Appendix G Extended Training Dynamics
We extend the training of 30B models from 80 steps to 120 steps with the Web environment. As shown in Figure 6, the results align with our analysis in Section 4.3. While both methods’ performance gradually converges, E-GRPO maintains a consistent lead over GRPO. Crucially, E-GRPO also continues to use fewer steps on average. This demonstrates E-GRPO’s dual advantages in both effectiveness and efficiency.
Appendix H Robust Matching Analysis
While exact string matching is computationally efficient and effectively mitigates reward hacking (see Appendix B.2), its rigidity presents practical challenges when faced with natural language variations such as alternative spellings, synonyms, or abbreviations. To address this limitation, we introduce a “Robust Matching” strategy. This approach involves first prompting an advanced LLM to generate a set of 5-10 plausible variations for each ground-truth entity, as detailed in the prompt below. Subsequently, we perform exact string matching against this expanded set of candidates. An entity is matched if any of its generated variations are found in the agent’s thought process.
With this “Robust Matching” strategy, we conduct the analysis of entity matching and correctness. Figure 7 visualizes the efficacy of this strategy. Compared with exact string matching, the “Robust Matching” strategy demonstrates a much stronger correlation with factual correctness. Specifically, for correct samples, the density peak at a match rate of 1.0 is significantly higher. Conversely, for incorrect samples, the peak at 0.0 is reduced, and the distribution becomes more spread out, better capturing “near-misses”.
This analysis confirms that “Robust Matching” serves as a more accurate proxy for factual correctness. Motivated by this finding, we conducted an additional RL experiment using the “Robust Matching” strategy, keeping all other settings identical to Web-30B-E-GRPO. The results are presented in Table 4.
| Strategy | GAIA | BrowseComp | BrowseComp-ZH | xbench-DS |
| Exact String Matching | 48.5 | 12.9 | 26.4 | 46.7 |
| Robust Matching | 49.2 | 12.9 | 26.8 | 47.0 |
The results are promising. The agent trained with “Robust Matching” achieves slight performance gains on three of the four deep research benchmarks, further validating the benefits of it. We consider the continued refinement of this strategy a promising direction for future work.
Appendix I Explanation of the Analysis Figure
Here we clarify the upper-right subfigure of Figure 1. For a question , we sample 8 responses for it. Each response matches entities. We calculate the average matched entities of correct responses and that of incorrect ones. For example, if only are correct, then correct rate is and wrong rate is . We compare these two rates for each question. The analysis shows that for most of the questions, correct rate is larger than wrong rate, i.e., correct responses of each question often match more entities than wrong ones. This establishes a strong correlation between the entity match rate and the correctness of the final answer.
Appendix J Use of Large Language Models (LLMs)
In this paper, we leverage Gemini-2.5-Pro (Google Team, 2025a) to polish paper writing grounded on our human-written content. The usage is limited to language polishing. We are responsible for any contributions generated by LLMs.